Performance Gloves, Inc. produces three sizes of sports gloves: small, medium, and large. A glove pattern is first stenciled onto leather in the Pattern Department. The stenciled patterns are then sent to the Cut and Sew Department, where the glove is cut and sewed together. Performance Gloves uses the multiple production department factory overhead rate method of allocating factory overhead costs. Its factory overhead costs were budgeted as follows: Pattern Department overhead $180,000 Cut and Sew Department overhead 297,600 Total $477,600 The direct labor estimated for each production department was as follows: Pattern Department 2,500 direct labor hours Cut and Sew Department 3,100 Total 5,600 direct labor hours Direct labor hours are used to allocate the production department overhead to the products. The direct labor hours per unit for each product for each production department were obtained from the engineering records as follows:

Production Departments Small Glove Medium Glove Large Glove
Pattern Department 0.05 0.06 0.07
Cut and Sew Department 0.07 0.09 0.11
Direct labor hours per unit 0.12 0.15 0.18

If required, round all per unit answers to the nearest cent.

a. Determine the two production department factory overhead rates.

Pattern Department $per dlh
Cut and Sew Department $per dlh

b. Use the two production department factory overhead rates to determine the factory overhead per unit for each product.

Small glove $ per unit
Medium glove $ per unit
Large glove $ per unit
Answer #1

A B с 1 a) 2 Pattern Department $ 72 per dlh 3 Cut and Sew Department $ 96 per dlh 4 5 b) 6 Small glove $ 10.32 per unit 7 Me

A B с per dih per dlh 1 a) 2 Pattern Department =180000/2500 3 Cut and Sew Department |=297600/3100 4 5 b) 6 Small glove =+(B